OpenStudy (anonymous):

The side lengths of a square are 7 in. What is the length of the diagonal? Question 13 options: 7 14 72√ 73√

OpenStudy (anonymous):

length of the diagonal = sqrt(2) * side of square

OpenStudy (anonymous):

this isnt the answer

OpenStudy (anonymous):

some problem with the third and fourth option given answer should be \[7\sqrt{2}\]

OpenStudy (anonymous):

@jhonyy9 it is a diagonal of a cube. Isn't it?

jhonyy9 (jhonyy9):

yes so than ?

OpenStudy (anonymous):

is it relevant to mention cube here?

jhonyy9 (jhonyy9):

so this is clearly sure that every sides are the same equale but for yo can calculi the diagonale of the square you need calculi first of all one diagonal of a side yes ?

OpenStudy (anonymous):

@jhonyy9 You are a champion . I assume you know better method than mine. So go for it, mate

OpenStudy (anonymous):

 square ABCD property of square - 1) sides of square are equal in length 2) all sides are perpendicular to each other i used pythagoras theorem to determine the length of diagonal of a square BD^2 = (7)^2 + (7)^2 BD = sqrt(2*49) = 7 sqrt(2) PS- square is a 2-dimensional figure and cube is 3-dimensional figure